iron 2 sulfate heptahydrate, FeSO4 + 7H2O is often used to treat patients with anemia due to iron mineral deficiency.


a) what is the molar mass of iron 2 sulfate heptahydrate?

b) calculate the percentage of iron in iron 2 sulfate heptahydrate.


1
Expert's answer
2021-03-25T04:16:45-0400

(a) Molar mass of "FeSO_4" is "278\dfrac{gm}{mol}"


(b)the percentage of iron in iron 2 sulfate heptahydrate


molar mass of Fe is "55.85\dfrac{gm}{mol}"


%percentage content ="\dfrac{55.85}{278}\times100=20.08" %

"%"